# Deep Transfer Learning: Model Framework and Error Analysis

### v1

@@ Appendix D Theorem cited during our proof
Theorem 5 (Theorem 3.2 in [83]).
@@
Theorem 6 (Lemma 2.3 in [83]).
@@
Datasets. For the classification datasets. VLCS [57] is a challenging dataset consists of 10,700 images. It has VOC2007 (V), LabelMe (L), Caltech (C) and Sun (S) four domains and 5 classes. PACS [58] is a dataset consisting of 9,900 images, which has 7 classes and 4 domains. TerraIncognita [59] is a geographic dataset consisting of images covering a diverse range of environments and objects. It has 4 domains and 24,700 images. OfficeHome [60] contains images collected from four distinct domains: Art (A), Clipart (C), Product (P), and Real-world (R) and corresponding annotations for classification tasks in office environments . It consists of 15,500 images divided into 65 object categories.
For the regression datasets. The tool wear dataset is from the 2010 PHM Challenge [61]. Only three data points, namely C1, C4, and C6, are labeled in the original dataset. The original dataset comprises sensing signals sampled at 50 kHz, encompassing cutting force signals, vibration signals in the x, y, and z dimensions, and acoustic emission signals during processing. The model employs cutting force signals in three directions as input and the average wear values of the three blades for each tool as output to establish transfer tasks.
@@
Hyperparameter search. For the classification task, we maintained a fixed batch size of 64 and utilized the Adam[84] optimizer with the default setting with the only different in learning rate. Both upstream and downstream training consisted of 50 epochs. The upstream training employed a constant learning rate of 0.00001 with the Adam optimizer. For the downstream task, the learning rate for the fully connected layers was set at 0.1, and using the CosineAnnealingLR scheduler. Also, the learning rate for EfficientNet’s feature extractor, which is the Q network, is 0.00001. The hyperparameter search focused solely on the penalty parameters, with values for $\lambda$ selected from {5, 10, 15}, $\tau$ from {0, 10} and $\mu$ from {0, 10}. The hyperparameter for the independence penalty term in the downstream task, $\kappa$ is from {5, 10, 15}.
@@
For the regression task, we maintained a fixed batch size of 50 and employed the Adam optimizer with settings identical to those used in the classification task, except for a variation in the learning rate. The upstream training consisted of 50 epochs and downstream training consisted of 100 epochs. The upstream training employed a learning rate of 0.2. For the downstream task, the learning rate for the fully connected layers was set at 0.02. Both training processes using the CosineAnnealingLR scheduler. The hyperparameter search focused solely on the penalty parameters, with values for $\lambda$ selected from {50, 100, 150, 200}, $\tau$ from {0, 10} and $\mu$ from {0, 10}. The parameters $\kappa$ for the independence penalty term in the downstream task is from {5, 10, 20}.
The size of our fully connected layer for dimensionality reduction is 2048*1024 in classification task, which means the shared feature dimensions in our experiment is 1024. Due to constraints in computational resources and time, we did not explicitly explore the impact of shared feature dimensions $r$ on downstream task performance. This applies to the parameters $\chi$, $\zeta$, and $d^{*}$ as well.
Results of regression dataset. For the improvement in regression dataset is not significant for every domain, so we just list the result here. The phrase ”c4, c6 -c1” in Table 4 represents using the data from domains c4 and c6 as upstream data, and the data from domain c1 as downstream data. The subsequent ”c1, c6 -c4” and ”c1, c4 -c6” have similar meanings.
Table 4: Fine-tuning accuracy on regression dataset and their domains
Transfer task/Algorithm
c4,c6 -c1 MAE RMSE R2
ERM 9.45 13.04 0.71
Ours 3.26 3.91 0.97
c1,c6 -c4 MAE RMSE R2
ERM 4.70 5.68 0.96
Ours 4.65 5.56 0.96
c1,c4 -c6 MAE RMSE R2
ERM 4.31 5.26 0.98
Ours 3.48 4.44 0.98
@@
Details about comparison with other works Due to time constraints and issues with the code logic, we aligned all training details as closely as possible except for the data input method. In the DomainBed codebase[67], the input data is structured such that each remaining domain inputs one batch size of samples, while our input method shuffles all remaining domains and inputs one batch size of samples. Therefore, to align as closely as possible, we set the batch size for upstream training to 21 and the batch size for downstream training to 64 when conducting comparative experiments in the DomainBed codebase. The typical hyperparameters for the compared algorithms are set to their default values in the DomainBed codebase.
@@
Error bar. Due to limited computational resources, we conducted downstream training three times using the checkpoints from the initial upstream training. For comparative results with other methods, particularly those using the DomainBed code [67], we repeated the tests by only varying the seed. The error bar represents the standard deviation of the best result from these experimental runs.
@@
Appendix F Limitations.
In all the theorems presented in this paper, we analyze the properties of the optimal empirical solutions obtained from our training procedure. However, in practice, it is challenging to ensure that training achieves the optimal empirical solution for most real-world problems solved using neural networks, and this is not unique to our method. Furthermore, in the celebrated paper [85], they give a result about wide (over-parameterized) neural networks satisfy the $PL^{*}$ condition, which explains the (S)GD convergence to a global minimum. Given the large number of parameters in our neural network model (ResNet50 with 25.6M parameters) compared to the sample size ($<$30k), this also implies that our training approach is somehow feasible.

### v3

@@ Appendix D Theorem cited during our proof
Theorem 5 (Theorem 3.2 in Jiao, Wang and Yang (2023)).
@@
By $\mathcal{SNN}_{d,1}(W,L,K)\subseteq\mathcal{NN}_{d,1}(W,L,K)\subseteq\mathcal{SNN}_{d,1}(W+1,L,K)$ (Proposition 2.1 in Jiao, Wang and Yang (2023)), the subsequent theorem can be employed during the proof of Theorem 1 and Theorem 3.
Theorem 6 (Lemma 2.3 in Jiao, Wang and Yang (2023)).
@@
Hyperparameter search. For the classification task, we maintained a fixed batch size of 64 and utilized the Adam (Kingma and Ba, 2015) optimizer with the default setting with the only different in learning rate. Both upstream and downstream training consisted of 50 epochs. The upstream training employed a constant learning rate of 0.00001 with the Adam optimizer. For the downstream task, the learning rate for the fully connected layers was set at 0.1, and using the CosineAnnealingLR scheduler. Also, the learning rate for EfficientNet’s feature extractor, which is the Q network, is 0.00001. The hyperparameter search focused solely on the penalty parameters, with values for $\lambda$ selected from {5, 10, 15}, $\tau$ from {0, 10} and $\mu$ from {0, 10}. The hyperparameter for the independence penalty term in the downstream task, $\kappa$ is from {5, 10, 15}.
@@
For the regression task, we maintained a fixed batch size of 50 and employed the Adam optimizer with settings identical to those used in the classification task, except for a variation in the learning rate. The upstream training consisted of 50 epochs and downstream training consisted of 100 epochs. The upstream training employed a learning rate of 0.2. For the downstream task, the learning rate for the fully connected layers was set at 0.005. Both training processes using the CosineAnnealingLR scheduler. The hyperparameter search focused solely on the penalty parameters, with values for $\lambda$ selected from {0.5, 1.0, 1.5, 2.0}, $\tau$ from {0, 0.2} and $\mu$ from {0, 0.5}. The parameter $\kappa$ for the independence penalty term in the downstream task is selected from $\{0.0625,0.125,0.25\}$, and $\chi$ for the $\ell_{1}$ penalty term is selected from $\{0.03125,0.0625\}$. To prevent certain components of the penalty terms from disproportionately affecting gradient calculations due to their larger values compared to other penalty terms, and to facilitate parameter selection, the penalty terms are normalized by their detached versions. The total loss is then computed as a weighted sum of the individual penalty terms, as follows: $\text{penalty}_{1}/\text{penalty}_{1}.detach()+\lambda*\text{penalty}_{2}/\text{penalty}_{2}.detach()+\tau*\text{penalty}_{2}/\text{penalty}_{2}.detach()+\mu*\text{penalty}_{3}/\text{penalty}_{3}.detach()$. Due to constraints in computational resources and time, we did not explicitly explore the impact of $\zeta$ and $d^{*}$ on downstream task performance.
@@
Details about comparison with other works. Due to time constraints and issues with the code logic, we aligned all training details as closely as possible except for the data input method. In the DomainBed codebase (Gulrajani and Lopez-Paz, 2020), the input data is structured such that each remaining domain inputs one batch size of samples, while our input method shuffles all remaining domains and inputs one batch size of samples. Therefore, to align as closely as possible, we set the batch size for upstream training to 21 and the batch size for downstream training to 64 when conducting comparative experiments in the DomainBed codebase. The typical hyperparameters for the compared algorithms are set to their default values in the DomainBed codebase.
[…]
Details of experiments exploring the effect of representation dimensionality To highlight the impact of sample size on estimation error, we trained the downstream models on OfficeHome using only 15% of the original training data. To investigate the influence of learning functions of different dimensions for downstream task model (8), we added three extra layers to the downstream network. Similarly, for the Wear dataset, we added two extra layers. For both datasets, we used cross-validation to select the optimal $\ell_{1}$ regularization hyperparameter.
@@
Error bar. Due to limited computational resources, we conducted downstream training three times using the checkpoints from the initial upstream training. For comparative results with other methods, particularly those using the DomainBed code (Gulrajani and Lopez-Paz, 2020), we repeated the tests by only varying the seed. The error bar represents the standard deviation of the best result from these experimental runs.